lilcutie2:

math help 15/24 = 25/x What is x? *

lilcutie2:

@AZ

AZ:

cross multiply so for example a/b = c/d would become ad = bc

AZ:

15/24 = 25/x 15 * x = 24 * 25 can you solve for x now?

lilcutie2:

15*x=600

lilcutie2:

then divide both sides by 15 and get x=40

lilcutie2:

@AZ am I right

lilcutie2:

@jhonyy9

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lilcutie2 @AZ am I right \(\color{#0cbb34}{\text{End of Quote}}\) That's correct! Good job

lilcutie2:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @lilcutie2 @AZ am I right \(\color{#0cbb34}{\text{End of Quote}}\) That's correct! Good job \(\color{#0cbb34}{\text{End of Quote}}\) how do we know if something is a proportion

AZ:

when you have two ratios or fractions equal to each other

AZ:

so if I say to make 1 cake, I need 3 eggs to make 3 cakes, how many eggs?? so the first fraction is 1 / 3 the second one is 3 / x so we can set them equal to make a proportion 1/3 = 3/x and we solve for x

lilcutie2:

Is 28/46 = 42/69 a proportion? * yes No I think no but I'm not sure

AZ:

oh like that

AZ:

so to make sure, you cross multiply just like we did earlier

AZ:

\(\dfrac{a}{b} = \dfrac{c}{d}\) \( ad = bc\)

lilcutie2:

so 28*42 and 46*69

AZ:

no no the numerator of the first fraction and the denominator of the other fraction

AZ:

and if the two numbers are equal, then it's a proportion

lilcutie2:

s0 28*69 and 42*46

lilcutie2:

they are a proportion

AZ:

Yup!